# FastTrackTr:Towards Real-time Multi-Object Tracking with Transformers

### v3

@@ 3.3 Historical Decoder and Encoder
Historical Decoder: Previous DETR-based MOT models [45, 22, 11] typically feed the final output queries of the previous frame, after some processing, directly into the decoder of the current frame to achieve tracking. This is an explicit tracking method, but it increases the number of queries, thereby increasing computational costs. To more efficiently utilize these historical queries, we replaced the self-attention mechanism in DETR’s decoder with a cross-attention mechanism to process the query features between consecutive frames. This significantly reduces the computational load. The differences between our decoder and the standard decoder are shown in Figure 3.
@@
Figure 3: Schematic of the flow of our historical-cross decoder and a general decoder. Compared to a general decoder, ours replaces self-attention with historical cross-attention.
@@
The workflow of the historical encoder is illustrated in Figure 4. After generating the mask, we fuse the decoder output $q_{f}^{t}$ with the historical information $q_{f}^{t-1^{\prime}}$ through an attention mechanism, computed as follows:
@@
where $q_{pos}$ represents a set of learnable positional encodings and the attention weights are computed via scaled dot-product as previously used in the decoder, followed by masking to suppress less relevant historical elements. Finally, the fused features are then processed by a feed-forward network to obtain $q^{t^{\prime}}_{f}$. For $t=1$, we initialize $q^{t-1^{\prime}}_{f}$ as $q^{t}_{f}$ due to the absence of prior history.
@@ 3.4 Matching and Association
Of course, considering that there are now many optimized methods for this process, we have also implemented several validated, simple, and effective techniques to enhance this workflow. Specific improvements and their impacts can be found in the appendix.
@@ 4 Experiments
Table 1: Overall Latency comparison between FastTrackTr and existing methods on the Dancetrack [34] val. * denotes that these methods employ the same YOLOX [13] detector, and the latency of this detector is accounted for in this table. Similarly, the latency of our method accounts for its associated time consumption, and these are highlighted in color. $\text{Dec}^{3}$ refers to a 3-layer decoder, while R18 refers to ResNet18. Unless otherwise specified, all Transformer-based models have ResNet50 for Backbone and 6 layers for decoder.
Methods Image Res. Numerical Precision Overall Latency(ms) HOTA Overall Latency with TensorRT FP16(ms)
Deep OC-SORT*[21] 1333×800 FP16 36.6ms (27.8fps) 58.5 13.5ms(74.1fps)
SUSHI*[6] 1333×800 FP32/FP16 42.7ms (23.4fps) / 38.0ms (26.3fps) 58.7 -
PuTR*[16] 1333×800 FP32/FP16 29.9ms (33.4fps) / 27.7ms (36.1fps) 54.5 -
@@ 4.1 Setting
Datasets: Our proposed method mainly evaluated on DanceTrack [34] and SportsMOT [7], two recently introduced datasets with large-scale training data, which helps prevent overfitting. Both datasets also provide official validation sets, allowing for exploratory experiments. Additionally, we present results on the MOT17 [23] benchmark.
Metrics: We assessed our method using widely recognized MOT evaluation metrics. The primary metrics employed for evaluation included HOTA, AssA, DetA ,IDF1 and MOTA [28, 20]. These metrics collectively provided a comprehensive and robust evaluation of our method’s tracking performance.
Implementation Details: For inference speed, our FastTrackTr is built on RT-DETR [50], with ResNet50 [14] as the backbone unless specified otherwise. The model is implemented in PyTorch [25], and the training is conducted on 8 NVIDIA RTX 4090 GPUs. The sequence length for training is set to 8 unless otherwise noted. All ablation experiments use an image size of $640\times 640$. For further testing of the best tracking accuracy of our method, we use an image size of $800\times 1333$ when comparing with SOTA methods. The Historical Encoder during training applies random mask transformations, with the probability related to the number of ground-truth objects. Given the number of ground-truth objects $n_{gt}$ and the number of queries $n_{q}$, the probability of a non-masked flip is $\min(0.5,2\cdot n_{gt}/n_{q})$, while for the rest, the probability is $\min(\max(0.1,n_{gt}/n_{q}),0.2)$. For all three datasets, $n_{q}$ is set to 300. Dancetrack and SportMOT are trained for 27 epochs, while MOT17 is trained for 80 epochs. The optimizer is AdwmW [19] with a learning rate of $10^{-4}$ and weight decay of $5.0\times 10^{-4}$. Due to GPU memory limitations, the batch size is set to 4 for smaller images and 1 for larger images. Lastly, for all datasets, the final output dimension of ID embedding is 256. All models are implemented in PyTorch by default unless explicitly stated otherwise. The complete codebase remains exclusively Python throughout our experiments.
@@ 4.2 Main Result
Comprehensive Comparison of Speed and Accuracy: In RT-DETR, a standardized benchmark has been established for real-time detectors. Within this benchmark, all models undergo acceleration via TensorRT, with computational latency measurements exclusively encompassing inference and post-processing durations. This paper adopts the same configuration while substituting the detection dataset COCO2017 [15] with the MOT dataset Dancetrack . However, this substitution reveals a critical implementation challenge: most transformer-based MOT methods [33, 22, 45, 41, 11, 44, 12, 16] developed in PyTorch incorporate dynamic computational graphs during inference, presumably to reduce computational overhead or simplify implementation. These dynamic elements create significant compatibility issues (specifically non-fixed tensor shapes and conditional branching) when converting models to TensorRT - challenges that remain unresolved in our current implementation. To ensure fair comparison, Table 1 comprehensively presents both speed and accuracy metrics across tracking methods using various numerical precisions in PyTorch. Notably, for our proposed models and the TBD method, only the network benefits from TensorRT acceleration, while the association module retains its original Python implementation.
@@
Comparison of State-Of-The-Art: To better demonstrate the superiority of our FastTrackTr, we employed nearly all available data augmentations and used larger image sizes. Tables 2, 4, and 3 respectively showcase the maximum tracking performance of FastTrackTr on DanceTrack, SportsMOT, and MOT17 benchmark. Similar to many previous works [51, 49, 11, 45], we utilized the CrowdHuman [30] dataset for data augmentation on MOT17 and DanceTrack. It is important to note that although our method still does not outperform CNN-based models on the MOT17 benchmark, the underlying reason remains the well-discussed issue: for transformer models, the amount of data in the MOT17 dataset is too limited. Nevertheless, our method has surpassed all previous transformer-based approaches. This indicates that our method performs better even when the dataset images do not allow for sufficient training of the model, suggesting that our approach has potential for practical application in real-world settings, rather than being confined mostly to academic research like previous transformer methods, with few exceptions.
@@ 4.3 Ablation Study
The ablation study in Table 5 highlights the progressive enhancements of FastTrackTr on the DanceTrack validation set. Starting with a baseline HOTA of 47.3, adding the Decoder increases HOTA to 51.6, AssA to 37.2, and IDF1 to 51.3. Incorporating both Decoder and Encoder further boosts HOTA to 53.8, AssA to 39.9, and IDF1 to 53.5. Finally, including the Mask module achieves the highest scores with HOTA of 54.1, AssA of 40.0, and IDF1 of 54.7. These results demonstrate that each component significantly contributes to the overall tracking performance of FastTrackTr.
@@
In Table 6, we demonstrate the impact of different trajectory cross-attention mechanisms within the trajectory decoder on tracking performance. Here, $q^{t}$ represents the original query, while $q_{f}^{t-1}$ denotes the output of the decoder from the previous frame, with ’concat’ indicating their concatenation. It is important to note that in this table, $q_{f}^{t-1}$ has not passed through the historical encoder, hence its performance is not as good as that of the final model.
@@
Table 7 analyzes the effect of different length of video clips during training. Utilizing 8 frames yields the highest HOTA, demonstrating superior tracking accuracy. Compared to using fewer frames, 8 frames offer substantial performance improvements while maintaining efficient training times. Beyond 8 frames, metrics show minimal or negligible gains, indicating diminishing returns with additional frames. Thus, selecting 8 training frames provides an optimal balance between enhanced performance and training efficiency for FastTrackTr.
Table 7: The length of video clips during training.
Training frames HOTA DetA AssA IDF1
3 51.7 72.8 36.7 50.9
5 53.5 74.2 38.8 53.5
8 54.1 75.6 40.0 54.7
9 54.0 75.4 39.6 54.8
10 54.1 75.8 39.8 54.6
In Table 8, we demonstrate the impact of using circle loss [35], triplet loss [29], and one-hot encoding loss of the JDE [37] on our method performance during the training of ID embeddings. From the result, circle loss consistently outperforms both triplet loss and one-hot encoding loss across all metrics. Specifically, circle loss achieves the highest tracking performance. In comparison, triplet loss shows the lowest performance, and one-hot encoding loss provides intermediate results. These results demonstrate that circle loss effectively enhances the training of ID embeddings, leading to superior tracking performance in our method.
Table 8: The impact of various ID Embedings losses on performance.
@@
5 Conclusion and Limitations

### v4

@@ 3.3 Historical Decoder and Encoder
While preceding sections have detailed the construction of an efficient Transformer-based MOT framework, the inherent temporal aggregation capabilities of Transformers remain underexploited in our current implementation. This section introduces a systematic methodology for efficiently leveraging historical information within the current frame.
Historical Decoder: Previous DETR-based MOT models [9, 8, 11] typically feed the final output queries of the previous frame, after some processing, directly into the decoder of the current frame to achieve tracking. This is an explicit tracking method, but it increases the number of queries, thereby increasing computational costs. To more efficiently utilize these historical queries, we replaced the self-attention mechanism in DETR’s decoder with a cross-attention mechanism to process the query features between consecutive frames. This significantly reduces the computational load. The differences between our decoder and the standard decoder are shown in Fig. 3.
@@
Figure 3: Schematic of the flow of our historical-cross decoder and a general decoder. Compared to a general decoder, ours replaces self-attention with historical cross-attention. In our historical cross-attention mechanism, the setting of $q$ remains the same, while $k$ and $v$ incorporate information from historical frames.
@@
The workflow of the historical encoder is illustrated in Fig. 4. After generating the mask, we fuse the decoder output $q_{f}^{t}$ with the historical information $q_{f}^{t-1^{\prime}}$ through an attention mechanism, computed as follows:
@@
where $q_{pos}$ represents positional encodings and the attention weights are computed via scaled dot-product as previously used in the decoder, followed by masking to suppress less relevant historical elements. Finally, the fused features are then processed by a feed-forward network to obtain $q^{t^{\prime}}_{f}$. For $t=1$, we initialize $q^{t-1^{\prime}}_{f}$ as $q^{t}_{f}$ due to the absence of prior history.
@@ 3.4 Matching and Association
Meanwhile, to further improve the model’s performance, similar to MMTrack [5], we adopted the iterative strategy of BYTE [31] and ID features, which is consistent with the proposals of [22, 16].
@@ 4 Experiments
Table 1: Overall Latency comparison between FastTrackTr and existing methods on the Dancetrack [24] val. * denotes that these methods employ the same YOLOX [34] detector, and the latency of this detector is accounted for in this table. Similarly, the latency of our method accounts for its associated time consumption, and these are highlighted in color. Unless otherwise specified, all Transformer-based models have ResNet50 for Backbone and 6 layers for decoder besides PuTR as it is a method of TBD paradigm. Not providing methods to accelerate TensorRT implies that they may not be well-suited for it.
Methods Image Res. Numerical Precision Overall Latency(ms) HOTA
Overall Latency with
TensorRT FP16(ms)
DiffMOT*[35] 1333×800 FP16 43.3ms (23.1fps) 55.7 22.3ms(44.8fps)
PuTR*[30] 1333×800 FP32/FP16 29.9ms (33.4fps) / 27.7ms (36.1fps) 54.5 15.8ms(63.3fps)
@@ 4.1 Setting
Datasets: To evaluate the tracking performance of FastTrackTr in real-world scenarios, we selected five datasets for testing: Dancetrack [24], SportMOT [25], MOT17 [26], BDD100k [28], and VisDrone2019 [27]. These datasets comprehensively encompass most real-world scenarios—for instance, MOT17 for surveillance, BDD100k for driving, VisDrone2019 for drone-based applications, and DanceTrack/SportMOT for sports-related contexts. Finally, we also tested our model on our own physical platform.
Metrics: We assessed our method using widely recognized MOT evaluation metrics. The primary metrics employed for evaluation included HOTA, AssA, DetA ,IDF1 and MOTA [36, 37]. These metrics collectively provided a comprehensive and robust evaluation of our method’s tracking performance. We also present additional evaluation metrics, including false positives (FP), false negatives (FN), identity switches (IDS), detection accuracy (DetA), association accuracy (AssA), and frame rate (FPS). And, for BDD100K, to better evaluate the performance of multiclass and multiobject tracking, we use tracking accuracy of all things (TETA) [28], localization accuracy (LocA), association accuracy (AssocA) and classification accuracy (ClsA) metrics.
Implementation Details: For inference speed, our FastTrackTr is built on RT-DETR [19], with ResNet50 as the backbone unless specified otherwise. The model is implemented in PyTorch, and the training is conducted on 8 NVIDIA RTX 4090 GPUs. The sequence length for training is set to 8 unless otherwise noted. All ablation experiments use an image size of $\times 640$ 640. For further testing of the best tracking accuracy of our method, we use an image size of 800×1333800$\times$ 1333800 × 1333 when comparing with SOTA methods. The Historical Encoder during training applies random mask transformations, with the probability related to the number of ground-truth objects. Given the number of ground-truth objects $n_{gt}$ and the number of queries $n_{q}$, the probability of a non-masked flip is $\min(0.5,2\cdot n_{gt}/n_{q})$, while for the rest, the probability is $\min(\max(0.1,n_{gt}/n_{q}),0.2)$. For all three datasets, $n_{q}$ is set to 300. Dancetrack, BDD100k and SportMOT are trained for 27 epochs, while MOT17 and VisDrone2019 are trained for 80 epochs. The optimizer is AdwmW with a learning rate of $10^{-4}$ and weight decay of $5.0\times 10^{-4}$. Due to GPU memory limitations, the batch size is set to 4 for smaller images and 1 for larger images on one GPU. Lastly, for all datasets, the final output dimension of ID embedding is 256. All models are implemented in PyTorch by default unless explicitly stated otherwise. The complete codebase remains exclusively Python throughout our experiments.
@@ 4.2 Main Result
Comprehensive Comparison of Speed and Accuracy: In RT-DETR, a standardized benchmark has been established for real-time detectors. Within this benchmark, all models undergo acceleration via TensorRT, with computational latency measurements exclusively encompassing inference and post-processing durations. This paper adopts the same configuration while substituting the detection dataset with the MOT dataset Dancetrack . However, this substitution reveals a critical implementation challenge: many transformer-based MOT methods [7, 8, 9, 10, 11, 12, 13] developed in PyTorch incorporate dynamic computational graphs during inference, presumably to reduce computational overhead or simplify implementation. These dynamic elements create significant compatibility issues (specifically non-fixed tensor shapes and conditional branching) when converting models to TensorRT - challenges that remain unresolved in our current implementation. To ensure fair comparison, Table 1 comprehensively presents both speed and accuracy metrics across tracking methods using various numerical precisions in PyTorch. Notably, for FastTrackTr and the TBD method, only the network benefits from TensorRT acceleration, while the association module retains its original Python implementation and similar to real-time detection, in this paper, we only measure the time consumption of the tracking system, and the time for data preprocessing and post-processing is not included.
@@
Comparison of State-Of-The-Art: To better demonstrate the superiority of our FastTrackTr, we employed nearly all available data augmentations and used larger image sizes. Tables 2, 3 ,4, 5 and 6 respectively showcase the maximum tracking performance of FastTrackTr on all benchmarks. Similar to many previous works [32, 29, 11, 9], we utilized the CrowdHuman [39] dataset for data augmentation on MOT17 and DanceTrack. It is important to note that although our method still does not outperform CNN-based models on the MOT17 benchmark, the underlying reason remains the well-discussed issue: for transformer models, the amount of data in the MOT17 dataset is too limited. Nevertheless, our method has surpassed all previous transformer-based approaches. This indicates that our method performs better even when the dataset images do not allow for sufficient training of the model, suggesting that our approach has potential for practical application in real-world settings, rather than being confined mostly to academic research like previous transformer methods, with few exceptions. And, We also provide various datasets and visualization results verified on physical objects, as shown in Fig. 5.
@@
Table 5: Comparison With the SOTA Methods on the VisDrone2019 Test Set
Methods MOTA $\uparrow$ IDF1 $\uparrow$ FP $\downarrow$ FN $\downarrow$ IDS $\downarrow$
MOTR [9] 22.8 41.4 28407 147937 959
TrackFormer [8] 25.0 30.5 25856 141526 4840
MMTrack [5] 36.7 54. 7 23849 120839 545
FastTrackTr 42.8 62.4 16474 85367 309
Table 6: Comparison to existing methods on the BDD100K validation set. Best results are marked in bold.
Methods TETA LocA AssocA ClsA
QDTrack [38] 47.8 45.8 48.5 49.2
MOTRv2 [29] 54.9 49.5 51.9 63.1
CO-MOT [10] 52.8 48.7 56 2 63.6
FastTrackTr 55.1 49.8 54 2 65.3
Figure 5: The visualization of FastTrackTr on various datasets. Ours represents the actual measured data of our driverless car on the road, and the model used here is trained on BDD100k.
@@ 4.3 Ablation Study
The ablation study in Table 7 highlights the progressive enhancements of FastTrackTr on the DanceTrack validation set. Starting with a baseline HOTA of 47.3, adding the Decoder increases HOTA to 51.6. Incorporating both Decoder and Encoder further boosts HOTA to 53.8. Finally, including the Mask module achieves the highest scores with HOTA of 54.1. In terms of the improvement degree, the Decoder with the introduction of historical information brings about the most significant performance enhancement. Adding an Encoder that further aggregates the temporal information of historical frames also contributes to a performance increase. The introduction of mask not only shields off useless historical information but also the randomness generated during its training can boost the generalization ability of the model.
@@
In Table 8, we demonstrate the impact of different trajectory cross-attention mechanisms within the trajectory decoder on tracking performance. Here, $q^{t}$ represents the original query, while $q_{f}^{t-1}$ denotes the output of the decoder from the previous frame, with ’concat’ indicating their concatenation. It is important to note that in this table, $q_{f}^{t-1}$ has not passed through the historical encoder, hence its performance is not as good as that of the final model. The advantage of Concat lies in combining the features of the current query and historical states. By leveraging the complementarity of spatiotemporal information and the dynamic balancing of the attention mechanism, it significantly boosts the joint performance of detection and tracking.
@@
In Table 9, we demonstrate the impact of using circle loss [33], triplet loss [40], and one-hot encoding loss of the JDE [23] on our method performance during the training of ID embeddings. From the result, circle loss consistently outperforms both triplet loss and one-hot encoding loss across all metrics. Specifically, circle loss achieves the highest tracking performance. In comparison, triplet loss shows the lowest performance, and one-hot encoding loss provides intermediate results. These results demonstrate that circle loss effectively enhances the training of ID embeddings for, leading to superior tracking performance in our method.
Table 9: The impact of various ID Embeddings losses on performance.
@@
4.4 Experiments in Actual Application Scenarios
Figure 6: Unmanned vehicle for physical verification.
Table 10: The speed and accuracy of our method and some existing methods using TensorRT FP16 precision on edge devices on our own verification platform and the accuracy of these models on Dancetrack are provided for reference.
Methods Image Res. Overall Latency(ms) HOTA
Deep OC-SORT[14] 1333 $\times$ 800 163.9 ms(6.1 fps) 58.5
PuTR [30] 1333 $\times$ 800 67.3ms(14.8fps) 54.5
OC-SORT [16] 1333 $\times$ 800 54.3 ms(18.4 fps) 52.1
FastTrackTr 1333 $\times$ 800 48.7 ms(20.5 fps) 56.9
Deep OC-SORT [14] 640 $\times$ 640 141.1 ms(7.1 fps) -
OC-SORT [16] 640 $\times$ 640 36.7 ms(27.2 fps) -
FastTrackTr 640 $\times$ 640 35.6 ms(28.1 fps) 54.8
FastTrackTr-Dec3 640 $\times$ 640 30.1ms(33.2fps) 51.2
FastTrackTr-R18-Dec3 640 $\times$ 640 16.8ms(59.2fps) 47.4
To validate the performance of our method in practical industrial environments, we conducted tests on the autonomous vehicle shown in Fig. 6. The core processor of this autonomous vehicle is NVIDIA’s Jetson AGX Orin, and the vehicle is equipped with sensors such as a lidar, an IMU, a GPS module, a fisheye camera, and an RGB camera. A Redmi AX3000 router is used to build an in-vehicle local area network, satisfying the network communication between the host and the lidar as well as the data transmission between the client and the autonomous vehicle host. To further simulate real-world scenarios, in addition to the tracking method, a SLAM algorithm is also running on the autonomous vehicle. The final time consumption of each tracking method is shown in Table 10, and the indicators marked in the table are the accuracy of the methods on Dancetrack. Since the official details on how OC - SORT trains the YOLOX detector are not provided, the accuracy of our trained model is not ideal, which may cause some misunderstandings. Therefore, we do not provide the specific tracking accuracy of OC - SORT and Deep OC - SORT under the image size of 640×640, and only provide the speed.
5 Conclusion